OpenStudy (anonymous):

Which statement describes the translation of y = -1 over 5(x + 5)2 + 2 from standard position

OpenStudy (anonymous):

Forgot to put this are the things given a. Moved up and to the right. b. Moved up and to the left. c. Moved down and to the left. d. Moved down and to the right.

OpenStudy (dumbcow):

im assuming the parent function here would be \[y = \frac{1}{x^{2}+2}\] oh nevermind , misread the question ok \[y = a(x-h)^{2}+k\] if h >0, shift right ......h<0 , shift left k>0, shift up ............k<0, shift down

OpenStudy (anonymous):

So would that be a b c or d lol

OpenStudy (dumbcow):

thats for you to decide, im here to give you the tools to help look at function, is h positive or negative ? is k positive or negative?

OpenStudy (anonymous):

Both positive I think

OpenStudy (dumbcow):

almost...k is positive because its " +2" h is tricky...(x+5) = (x-(-5)) so really h = -5

OpenStudy (dumbcow):

general rule: take whats inside parenthesis and set equal to 0 --> x+5 = 0 --> x = -5 ......this is your h value

OpenStudy (anonymous):

So I am thinking it would be move up to the left

OpenStudy (dumbcow):

yep...im speaking in general terms so you can do any problem no matter what the numbers are

OpenStudy (anonymous):

So it would be move up and to the left?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (dumbcow):

welcome :)